# A systematic review of metaheuristics-based and machine learning-driven intrusion detection systems in IoT

### v3

@@ 5.2 RQ2: What are the most commonly used datasets and evaluation metrics for IoT-IDS assessment?
Our investigation reveals that the well-known NSL-KDD dataset [nsl_tavallaee2009detailed] is extensively used by IoT-IDSs. It consists of 148,517 records extracted from the 5,209,458 samples of the oldest benchmark dataset, KDDCup-99 [kdd_99] by removing redundant records. Both of these datasets contain 41 features and 5 target classes. In 2015, the UNSW-NB15 dataset [unsw_moustafa2015unsw] was created having 49 features and 10 target classes. Importantly, since this dataset does not have any outdated features of attacks, researchers tend to test their works on this. Additionally, the BoT-IoT [bot_koroniotis2019towards] and CICIDS-2017 [cic_sharafaldin2018toward] datasets are widely used as well. However, Table 7 states that CICIDS-2017 contains around 80 features and 8 classes; whereas BoT-IoT is a huge dataset with 73,360,900 records and 46 features and 7 classes. Several systems also utilize the TON-IoT [ton_moustafa2021new] and N-BaIoT [nbaiot_meidan2018n] datasets. TON-IoT is a large dataset with 22,339,021 records in total, of which 461,043 are dedicated to training and testing purposes. On the other hand, N-BaIoT consists of 23 features and 7,062,606 records. The proportion of each dataset used in the existing detection systems is shown in Figure 7. Apart from these mostly used datasets, the IoTID20 dataset [iotid20_ullah2020scheme] is used in 5 papers. Additionally, AWID [awid_kolias2015intrusion], WSN-DS [wsnds_almomani2016wsn], and RPL-NIDDS17 [rpl_verma2019evaluation] datasets are employed by three systems each; whereas, the rest of the works utilize distinct datasets.
@@
| NSL-KDD (2009) [nsl_tavallaee2009detailed] | Benign, Probe, DoS, U2R, R2L | 51.89 + 35.94 + 9.48 + 2.52 + 0.17 | 41 | 125,973 | 22,544 | 148,517 | ✓ |
| UNSW-NB15 (2015) [unsw_moustafa2015unsw] | Benign, Fuzzers, Analysis, Backdoors, DoS, Exploits, Generic, Reconnaissance, Shellcode, Worms | 87.36 + 0.95 + 0.11 + 0.09 + 0.64 + 1.75 + 8.48 + 0.55 + 0.06 + 0.01 | 49 | 175,341 | 82,332 | ✓ | |
| BoT-IoT (2019) [bot_koroniotis2019towards] | Benign, DoS, DDoS, Reconnaissance, Information Theft | 0.01 + 44.96 + 52.54 + 2.48 + 0.002 | 46 | - | - | 73,360,900 | ✓ |
| CICIDS-2017 (2017) [cic_sharafaldin2018toward] | Benign, DoS Hulk, PortScan, DDoS, DoS GoldenEye, FTP-Patator, SSH-Patator, DoS slowloris, DoS Slowhttptest, Bot, Web Attack – Brute Force, Web Attack – XSS, Infiltration, Web Attack – SQL Injection, Heartbleed | 83.34 + 8.16 + 5.61 + 1.48 + 0.36 + 0.28 + 0.21 + 0.20 + 0.19 + 0.07 + 0.05 + 0.02 + 0.00 + 0.00 + 0.00 | 80 | - | - | 2,830,743 | ✓ |
| KDDCup-99 (1999) [kdd_99] | Benign, Probe, DoS, U2R, R2L | 19.86 + 0.84 + 79.30 + 0.0 + 0.02 | 41 | 4,898,431 | 311,027 | 5,209,458 | ✓ |
| TON-IoT (2021) [ton_moustafa2021new] | Benign, Backdoor, DDoS, DoS, Injection, MITM, Password, Ransomware, Scanning, XSS | 3.56 + 2.27 + 27.60 + 15.11 + 2.03 + 0.00 + 7.69 + 0.33 + 31.96 + 9.44 | 46 | 461,043 | 22,339,021 | ✓ | |
| N-BaIoT (2018) [nbaiot_meidan2018n] | Benign, mirai_udp, gafgyt_udp, gafgyt_tcp, mirai_syn, mirai_ack, mirai_scan, mirai_udpplain, gafgyt_combo, gafgyt_junk, gafgyt_scan | 7.87 + 17.41 + 13.40 + 12.17 + 10.38 + 9.11 + 7.62 + 7.41 + 7.29 + 3.71 + 3.61 | 23 | - | - | 7,062,606 | mode-rately |
@@
5.3 RQ3: What are the relations between the optimization algorithms and classification methods with the datasets?
@@
5.3.1 Connecting dots among Metaheuristics, Datasets, and Machine Learning
@@
| NSL-KDD | [33_phalguna2021hybrid] | GWO-PSO | FS | RF | - | 99.97 | 99.96 | - | Q2 |
| [13_krishna2021attack] | LOA-FOA | FS | RF | - | 99.98 | 99.73 | AUC=99.76 | Q2 | |
| [56_xu2023iot] | BGWO | FS | EL (XGBoost) | - | 99.9427 | 99.9426 | - | Q2 | |
| [113_li2024cooperative] | BHO | FS | Parallel CNNs | - | 99.8928 | 99.89 | Q1 | ||
| [42_asgharzadeh2023anomaly] | BMECapSA | FS | CNN | 18 | 99.85 | 99.85 | FAR=0.0019, FNR=0.001 | Q1 | |
| [68_sandhya2021enhancing] | SMO | FS | RF | - | 99.675 | 99.9325 | AUC=99.3025 | Q2 | |
| [44_keserwani2021smart] | GWO-PSO | FS | RF | - | 99.66 | - | - | Q1 | |
| [67_gupta2022hybrid] | HCSGA | FS | DLHNN | - | 99.52 | 97.16 | - | Q1 | |
| [75_jothi2023wils] | WOA | PT | LSTM | N/A | 99.5 | - | Specificity=98.45 | Q1 | |
| [52_pan2021lightweight] | Compact SCA | HPT | KNN | N/A | 99.327 | - | FAR=0.5848 | Q2 | |
| [45_karthikeyan2024firefly] | GWO,FOA | PT, FS | SVM | - | 99.29 | 96.23 | FAR=99.59, AUC=98.51 | Q1 | |
| [90_biju2024evaluated] | EBSA | PT | DBN | N/A | 98.96 | 99.13 | - | Q2 | |
| [41_ethala2022hybrid] | SMO-HPSO | FS | RF | 22 | 98.98 | 98.59 | AUC=99.81 | Q1 | |
| [81_hanafi2024intrusion] | IBGJO | FS | LSTM | - | 98.93 | 98.17 | - | Q1 | |
| [110_maazalahi2024k] | ASO-EO, FOA | FS | k-means | - | 98.9 | 100 | - | Q1 | |
| [46_jayalatchumy2024improved] | enhanced CrSA | FS | EL | 11 | 99 | 98.14 | - | Q1 | |
| UNSW-NB15 | [8_savanovic2023intrusion] | modified FOA | HPT | EL(XGBoost) | N/A | 99.98 | 99.99 | AUC-ROC=1 | Q1 |
| [91_saheed2024voting] | GWO | HPT | EL | N/A | 100 | 99.745 | FAR=1.5, ROC=99.4 | Q1 | |
| [66_gadekallu2023moth] | MFO | FS | EL | 14 | 100 | 99.75 | - | Q1 | |
| [64_om2022effective] | CSA | PT | HKCAE | N/A | 99.7 | 98.9 | Specificity=98.3 | Q2 | |
| [20_alghamdi2022hybrid] | PO | PT | CFNN | N/A | 99.46 | 99.76 | - | Q2 | |
| [14_dey2023metaheuristic] | BGSA-BGWO | FS | EL | 4 | 99.41 | 99.33 | FAR=0.03 | Q2 | |
| [110_maazalahi2024k] | ASO-EO, FOA | FS | k-means | - | 99.1 | 99.4 | - | Q1 | |
| [75_jothi2023wils] | WOA | PT | LSTM | N/A | 99.1 | - | Specificity=98.99 | Q1 | |
| [80_bajpai2024hybrid] | WOA | HPT | GRU | N/A | 99 | - | - | Q2 | |
| [60_gangula2023intrusion] | FOA | FS | EL | - | 98.89 | 98.91 | AUC=99.79 | Q2 | |
| [88_rabie2024novel] | DRFO | FS | DBRBF | - | 98.5 | 98.5 | FAR=8.2 | Q1 | |
| [113_li2024cooperative] | BHO | FS | Parallel CNNs | - | 97.7217 | 97.56 | Q1 | ||
| [54_jeyaselvi2023highly] | EXPSO-STFA | FS | LAANN | - | 95.65 | 95.64 | Specificity=92.74, FNR=10.23, MCC=92.56 | Q1 | |
| BoT-IoT | [91_saheed2024voting] | GWO | HPT | EL | N/A | 99.98 | 99.955 | FAR=1.3, ROC=99.99 | Q1 |
| [64_om2022effective] | CSA | PT | HKCAE | N/A | 99.9 | 98.2 | Specificity=99.7 | Q2 | |
| [51_geetha2024cvs] | CVS | FS | FLN-ANN | - | 99.68 | 99.21 | Specificity=99.83 | Q1 | |
| [73_abd2023intrusion] | CSA | FS | CNN | - | 99.15 | 98.806 | - | Q1 | |
| [89_elsedimy2024novel] | IGWO | HPT | QSVM | N/A | 99.11 | 97.48 | - | Q1 | |
| [15_fatani2021iot] | TSO-DE | FS | CNN | - | 99.042 | 99.042 | FAR=0.00301 | Q1 | |
| [16_dahou2022intrusion] | RSA | FS | CNN | - | 99.02 | 99.07 | - | Q1 | |
| [49_fatani2021advanced] | AQUO | FS | CNN | 10 | 98.926 | 98.904 | - | Q1 | |
| [35_baniasadi2022novel] | NS-BPSO | PT | DCNN | N/A | 98.86 | - | Specificity=95.32, MSE=0.00053 | Q1 | |
| [74_saheed2023novel] | HAEMPSO | FS, HPT | DNN | - | 97.61 | - | - | Q1 | |
| [89_elsedimy2024novel] | IGWO | HPT | QSVM | N/A | 99.11 | 97.48 | - | Q1 | |
| [101_deore2022hybrid] | CCSO | PT | Deep LSTM | N/A | 96.71 | - | Specificity=91.985 | Q1 | |
| CICIDS-2017 | [15_fatani2021iot] | TSODE | FS | CNN | - | 99.93 | 99.93 | FAR=0.00009 | Q1 |
| [16_dahou2022intrusion] | RSA | FS | CNN | - | 99.911 | 99.888 | - | Q1 | |
| [49_fatani2021advanced] | AQUO | FS | CNN | - | 99.911 | 99.888 | - | Q1 | |
| [73_abd2023intrusion] | CSA | FS | CNN | - | 99.911 | 99.888 | - | Q1 | |
| [51_geetha2024cvs] | CVS | FS | FLN-ANN | - | 99.77 | 99.72 | Specificity=99.92 | Q1 | |
| [81_hanafi2024intrusion] | IBGJO | FS | LSTM | - | 99.75 | 98.81 | - | Q1 | |
| [94_aljehane2024golden] | GJOA,SSA | FS, HPT | A-BiLSTM | - | 99.69 | 98.92 | MCC=98.74 | Q1 | |
| [44_keserwani2021smart] | GWO-PSO | FS | RF | - | 99.66 | - | - | Q1 | |
| [114_alamro2023modeling] | ALO, FPA | FS, HPT | CNN+LSTM | - | 99.55 | 99.55 | AUC=99.55 | Q1 | |
| [105_10541183] | IGC-SA | FS | AE-DNN | - | 99.4 | 99.4 | - | PAIS24 | |
| [18_fatani2023enhancing] | MGO | FS | CNN | - | 99.22 | 99.218 | G-Mean=99.218 | Q1 | |
| [104_10317882] | IWD-BBO | FS | FNN | - | 98.2339 | 99.0865 | - | Q1 | |
| [81_hanafi2024intrusion] | IBGJO | FS | LSTM | - | 99.75 | 98.81 | - | Q1 | |
| KDDCup-99 | [18_fatani2023enhancing] | MGO | FS | CNN | - | 99.941 | 99.942 | G-Mean=99.942 | Q1 |
| [49_fatani2021advanced] | AQUO | FS | CNN | - | 99.919 | 89.987 | - | Q1 | |
| [73_abd2023intrusion] | CSA | FS | CNN | - | 99.917 | 89.988 | - | Q1 | |
| [62_alghanam2023improved] | LS-PIO | FS | EL | 15 | 99.82 | 97.23 | FAR=6.9, TPR=99.23, AUC=96.32 | Q1 | |
| [44_keserwani2021smart] | GWO-PSO | FS | RF | - | 99.66 | - | - | Q1 | |
| [110_maazalahi2024k] | ASO-EO | FS | k-means | - | 96.1 | 100 | - | Q1 | |
| [54_jeyaselvi2023highly] | EXPSO-STFA | FS | LAANN | - | 95.65 | 95.64 | Specificity=92.74, FAR=14.52, FNR=10.23, MCC=92.56 | Q1 | |
| [15_fatani2021iot] | TSODE | FS | CNN | - | 92.064 | 90.007 | FAR=0.01989 | Q1 | |
| [16_dahou2022intrusion] | RSA | FS | CNN | - | 92.04 | 89.985 | - | Q1 | |
| TON_IoT | [42_asgharzadeh2023anomaly] | BMECapSA | FS | CNN | 12 | 99.99 | 99.99 | FAR=0.0001, FNR=0.00002 | Q1 |
| [105_10541183] | SA | FS | AE-DNN | - | 99.888 | 99.875 | - | PAIS24 | |
| [82_kethineni2024intrusion] | WHO | PT | fused CNN-BiGRU | N/A | 99.71 | 99.05 | - | Q1 | |
| [106_escorcia2023sea] | STFA,SpSO | FS, HPT | DBN | - | 99.51 | - | Specificty=99.36, MCC=60.36 | Q1 | |
| [58_nazir2023novel] | TS | FS | EL | 13 | 99.5 | - | FAR=0.004 | Q1 | |
| N-BaIoT | [56_xu2023iot] | BGWO | FS | EL (XGBoost) | - | 99.9941 | 99.9941 | - | Q2 |
| [13_krishna2021attack] | LOA-FOA | FS | RF | - | 99.86 | 99.86 | - | Q2 | |
| [33_phalguna2021hybrid] | GWO-PSO | FS | RF | - | 99.86 | 99.86 | - | Q2 | |
| [21_alweshah2022intrusion] | SSO-SA | FS | KNN | - | 98.7 | 99.8 | - | Q2 | |
| [23_alweshah2023intrusion] | EPC | FS | KNN | - | 98.2 | 99.4 | - | Q1 | |
| [29_bahaa2022novel] | APSO-WOA | HPT | CNN | N/A | 94.54 | - | JCC= 0.9 | Q1 |
Takeaway. According to the correlation analysis, the existing best IoT-IDSs have achieved 99.97%99.99% accuracy and 99.95%99.99% f1-score with NSL-KDD, UNSW-NB15, BoT-IoT, TON-IoT, and N-BaIoT datasets. Whereas, the performance slightly reduces while experimenting on the CICIDS-2017 and KDDCup-99 datasets (99.93%99.94% accuracy and f1-score). Overall, the metaheuristics and ML-integrated detection systems are effective in the IoT environment, irrespective of all widely used datasets. Another important observation is that existing intrusion detection datasets are extremely imbalanced since significant discrepancy is observed in data distribution between majority and minority classes (see Table 7). Consequently, it is not difficult to get a remarkable accuracy from any ML classifier [somasundaram2016data], which can mislead researchers. Therefore, “accuracy” should not be considered a trustworthy performance metric in such scenarios. In contrast, the F1-score is widely accepted as a reliable metric since it reflects the harmonic balance between precision and recall [bej2021loras]. For this reason, almost all top-notch IoT-IDSs measure F1-score along with accuracy and others.
@@
Further investigation in ensemble learning classifiers reveals the efficiency of using both traditional and deep machine learning algorithms. A significant portion of these classifiers are based on classic and deep ML, such as KNN+SVM+LSTM+MLP [60_gangula2023intrusion], RF+DT+MLP+KNN [91_saheed2024voting], and SVM+KNN+RF+LSTM [46_jayalatchumy2024improved]. In addition, a combination of classic ML classifiers is also observed, especially RFs [58_nazir2023novel], LR+RF+XGBoost [66_gadekallu2023moth], DT+AdaBoost+RF [14_dey2023metaheuristic], and OC-SVM+IF+LOF [62_alghanam2023improved]. Interestingly, Latif et al. [112_latif2024dtl] leveraged a CNN-based bootstrap ensemble classifier (Generic CNN+Xception+Inception+InseptionResntV2+EffcientNetV2L). Regarding optimization techniques, there is a consistent trend of employing nature-inspired metaheuristics, specifically FOA [8_savanovic2023intrusion, 60_gangula2023intrusion], MFO [66_gadekallu2023moth], BGWO [56_xu2023iot], CrSA [46_jayalatchumy2024improved], GWO [91_saheed2024voting], and GA [112_latif2024dtl]. In addition to these, search-based TS [58_nazir2023novel] along with hybrid metaheuristics BGSA-BGWO [14_dey2023metaheuristic] and LS-PIO [62_alghanam2023improved] are efficient in IoT-IDS. Moreover, almost all metaheuristics are used to select features; except in [8_savanovic2023intrusion], where modified FOA tunes hyperparameters of the XGBoost classifier.
5.3.2 Delve Into the Parameter and Hyperparameter Tuning Application

### v4

@@ V-B RQ2: What are the most commonly used datasets and evaluation metrics for IoT-IDS assessment?
Our investigation reveals that the well-known NSL-KDD dataset [162] is extensively used by IoT-IDSs. It consists of 148,517 records extracted from the 5,209,458 samples of the oldest benchmark dataset, KDDCup-99 [95] by removing redundant records. Both of these datasets contain 41 features and 5 target classes. In 2015, the UNSW-NB15 dataset [125] was created having 49 features and 10 target classes. Importantly, since this dataset does not have any outdated features of attacks, researchers tend to test their works on this. Additionally, the BoT-IoT [102] and CICIDS-2017 [155] datasets are widely used as well. However, Table VII states that CICIDS-2017 contains around 80 features and 8 classes; whereas BoT-IoT is a huge dataset with 73,360,900 records and 46 features and 7 classes. Several systems also utilize the TON-IoT [126] and N-BaIoT [118] datasets. TON-IoT is a large dataset with 22,339,021 records in total, of which 461,043 are dedicated to training and testing purposes. On the other hand, N-BaIoT consists of 23 features and 7,062,606 records. The proportion of each dataset used in the existing detection systems is shown in Figure 7. Apart from these mostly used datasets, the IoTID20 dataset [164] is used in 5 papers. Additionally, AWID [101], WSN-DS [15], and RPL-NIDDS17 [168] datasets are employed by three systems each; whereas, the rest of the works utilize distinct datasets.
@@
| NSL-KDD (2009) [162] | Benign, Probe, DoS, U2R, R2L | 51.89 + 35.94 + 9.48 + 2.52 + 0.17 | 41 | 125,973 | 22,544 | 148,517 | ✓ |
| UNSW-NB15 (2015) [125] | Benign, Fuzzers, Analysis, Backdoors, DoS, Exploits, Generic, Reconnaissance, Shellcode, Worms | 87.36 + 0.95 + 0.11 + 0.09 + 0.64 + 1.75 + 8.48 + 0.55 + 0.06 + 0.01 | 49 | 175,341 | 82,332 | ✓ | |
| BoT-IoT (2019) [102] | Benign, DoS, DDoS, Reconnaissance, Information Theft | 0.01 + 44.96 + 52.54 + 2.48 + 0.002 | 46 | - | - | 73,360,900 | ✓ |
| CICIDS-2017 (2017) [155] | Benign, DoS Hulk, PortScan, DDoS, DoS GoldenEye, FTP-Patator, SSH-Patator, DoS slowloris, DoS Slowhttptest, Bot, Web Attack – Brute Force, Web Attack – XSS, Infiltration, Web Attack – SQL Injection, Heartbleed | 83.34 + 8.16 + 5.61 + 1.48 + 0.36 + 0.28 + 0.21 + 0.20 + 0.19 + 0.07 + 0.05 + 0.02 + 0.00 + 0.00 + 0.00 | 80 | - | - | 2,830,743 | ✓ |
| KDDCup-99 (1999) [95] | Benign, Probe, DoS, U2R, R2L | 19.86 + 0.84 + 79.30 + 0.0 + 0.02 | 41 | 4,898,431 | 311,027 | 5,209,458 | ✓ |
| TON-IoT (2021) [126] | Benign, Backdoor, DDoS, DoS, Injection, MITM, Password, Ransomware, Scanning, XSS | 3.56 + 2.27 + 27.60 + 15.11 + 2.03 + 0.00 + 7.69 + 0.33 + 31.96 + 9.44 | 46 | 461,043 | 22,339,021 | ✓ | |
| N-BaIoT (2018) [118] | Benign, mirai_udp, gafgyt_udp, gafgyt_tcp, mirai_syn, mirai_ack, mirai_scan, mirai_udpplain, gafgyt_combo, gafgyt_junk, gafgyt_scan | 7.87 + 17.41 + 13.40 + 12.17 + 10.38 + 9.11 + 7.62 + 7.41 + 7.29 + 3.71 + 3.61 | 23 | - | - | 7,062,606 | mode-rately |
@@
V-C RQ3: What are the relations between the optimization algorithms and classification methods with the datasets?
@@
V-C1 Connecting dots among Metaheuristics, Datasets, and Machine Learning
@@
| NSL-KDD | [132] | GWO-PSO | FS | RF | - | 99.97 | 99.96 | - | Q2 |
| [103] | LOA-FOA | FS | RF | - | 99.98 | 99.73 | AUC=99.76 | Q2 | |
| [173] | BGWO | FS | EL (XGBoost) | - | 99.9427 | 99.9426 | - | Q2 | |
| [108] | BHO | FS | Parallel CNNs | - | 99.8928 | 99.89 | Q1 | ||
| [24] | BMECapSA | FS | CNN | 18 | 99.85 | 99.85 | FAR=0.0019, FNR=0.001 | Q1 | |
| [149] | SMO | FS | RF | - | 99.675 | 99.9325 | AUC=99.3025 | Q2 | |
| [96] | GWO-PSO | FS | RF | - | 99.66 | - | - | Q1 | |
| [74] | HCSGA | FS | DLHNN | - | 99.52 | 97.16 | - | Q1 | |
| [88] | WOA | PT | LSTM | N/A | 99.5 | - | Specificity=98.45 | Q1 | |
| [131] | Compact SCA | HPT | KNN | N/A | 99.327 | - | FAR=0.5848 | Q2 | |
| [92] | GWO,FOA | PT, FS | SVM | - | 99.29 | 96.23 | FAR=99.59, AUC=98.51 | Q1 | |
| [34] | EBSA | PT | DBN | N/A | 98.96 | 99.13 | - | Q2 | |
| [57] | SMO-HPSO | FS | RF | 22 | 98.98 | 98.59 | AUC=99.81 | Q1 | |
| [77] | IBGJO | FS | LSTM | - | 98.93 | 98.17 | - | Q1 | |
| [113] | ASO-EO, FOA | FS | k-means | - | 98.9 | 100 | - | Q1 | |
| [85] | enhanced CrSA | FS | EL | 11 | 99 | 98.14 | - | Q1 | |
| UNSW-NB15 | [152] | modified FOA | HPT | EL(XGBoost) | N/A | 99.98 | 99.99 | AUC-ROC=1 | Q1 |
| [145] | GWO | HPT | EL | N/A | 100 | 99.745 | FAR=1.5, ROC=99.4 | Q1 | |
| [66] | MFO | FS | EL | 14 | 100 | 99.75 | - | Q1 | |
| [128] | CSA | PT | HKCAE | N/A | 99.7 | 98.9 | Specificity=98.3 | Q2 | |
| [9] | PO | PT | CFNN | N/A | 99.46 | 99.76 | - | Q2 | |
| [45] | BGSA-BGWO | FS | EL | 4 | 99.41 | 99.33 | FAR=0.03 | Q2 | |
| [113] | ASO-EO, FOA | FS | k-means | - | 99.1 | 99.4 | - | Q1 | |
| [88] | WOA | PT | LSTM | N/A | 99.1 | - | Specificity=98.99 | Q1 | |
| [27] | WOA | HPT | GRU | N/A | 99 | - | - | Q2 | |
| [67] | FOA | FS | EL | - | 98.89 | 98.91 | AUC=99.79 | Q2 | |
| [137] | DRFO | FS | DBRBF | - | 98.5 | 98.5 | FAR=8.2 | Q1 | |
| [108] | BHO | FS | Parallel CNNs | - | 97.7217 | 97.56 | Q1 | ||
| [87] | EXPSO-STFA | FS | LAANN | - | 95.65 | 95.64 | Specificity=92.74, FNR=10.23, MCC=92.56 | Q1 | |
| BoT-IoT | [145] | GWO | HPT | EL | N/A | 99.98 | 99.955 | FAR=1.3, ROC=99.99 | Q1 |
| [128] | CSA | PT | HKCAE | N/A | 99.9 | 98.2 | Specificity=99.7 | Q2 | |
| [68] | CVS | FS | FLN-ANN | - | 99.68 | 99.21 | Specificity=99.83 | Q1 | |
| [1] | CSA | FS | CNN | - | 99.15 | 98.806 | - | Q1 | |
| [55] | IGWO | HPT | QSVM | N/A | 99.11 | 97.48 | - | Q1 | |
| [58] | TSO-DE | FS | CNN | - | 99.042 | 99.042 | FAR=0.00301 | Q1 | |
| [39] | RSA | FS | CNN | - | 99.02 | 99.07 | - | Q1 | |
| [60] | AQUO | FS | CNN | 10 | 98.926 | 98.904 | - | Q1 | |
| [28] | NS-BPSO | PT | DCNN | N/A | 98.86 | - | Specificity=95.32, MSE=0.00053 | Q1 | |
| [146] | HAEMPSO | FS, HPT | DNN | - | 97.61 | - | - | Q1 | |
| [55] | IGWO | HPT | QSVM | N/A | 99.11 | 97.48 | - | Q1 | |
| [43] | CCSO | PT | Deep LSTM | N/A | 96.71 | - | Specificity=91.985 | Q1 | |
| CICIDS-2017 | [58] | TSODE | FS | CNN | - | 99.93 | 99.93 | FAR=0.00009 | Q1 |
| [39] | RSA | FS | CNN | - | 99.911 | 99.888 | - | Q1 | |
| [60] | AQUO | FS | CNN | - | 99.911 | 99.888 | - | Q1 | |
| [1] | CSA | FS | CNN | - | 99.911 | 99.888 | - | Q1 | |
| [68] | CVS | FS | FLN-ANN | - | 99.77 | 99.72 | Specificity=99.92 | Q1 | |
| [77] | IBGJO | FS | LSTM | - | 99.75 | 98.81 | - | Q1 | |
| [12] | GJOA,SSA | FS, HPT | A-BiLSTM | - | 99.69 | 98.92 | MCC=98.74 | Q1 | |
| [96] | GWO-PSO | FS | RF | - | 99.66 | - | - | Q1 | |
| [8] | ALO, FPA | FS, HPT | CNN+LSTM | - | 99.55 | 99.55 | AUC=99.55 | Q1 | |
| [36] | IGC-SA | FS | AE-DNN | - | 99.4 | 99.4 | - | PAIS24 | |
| [59] | MGO | FS | CNN | - | 99.22 | 99.218 | G-Mean=99.218 | Q1 | |
| [73] | IWD-BBO | FS | FNN | - | 98.2339 | 99.0865 | - | Q1 | |
| [77] | IBGJO | FS | LSTM | - | 99.75 | 98.81 | - | Q1 | |
| KDDCup-99 | [59] | MGO | FS | CNN | - | 99.941 | 99.942 | G-Mean=99.942 | Q1 |
| [60] | AQUO | FS | CNN | - | 99.919 | 89.987 | - | Q1 | |
| [1] | CSA | FS | CNN | - | 99.917 | 89.988 | - | Q1 | |
| [10] | LS-PIO | FS | EL | 15 | 99.82 | 97.23 | FAR=6.9, TPR=99.23, AUC=96.32 | Q1 | |
| [96] | GWO-PSO | FS | RF | - | 99.66 | - | - | Q1 | |
| [113] | ASO-EO | FS | k-means | - | 96.1 | 100 | - | Q1 | |
| [87] | EXPSO-STFA | FS | LAANN | - | 95.65 | 95.64 | Specificity=92.74, FAR=14.52, FNR=10.23, MCC=92.56 | Q1 | |
| [58] | TSODE | FS | CNN | - | 92.064 | 90.007 | FAR=0.01989 | Q1 | |
| [39] | RSA | FS | CNN | - | 92.04 | 89.985 | - | Q1 | |
| TON_IoT | [24] | BMECapSA | FS | CNN | 12 | 99.99 | 99.99 | FAR=0.0001, FNR=0.00002 | Q1 |
| [36] | SA | FS | AE-DNN | - | 99.888 | 99.875 | - | PAIS24 | |
| [97] | WHO | PT | fused CNN-BiGRU | N/A | 99.71 | 99.05 | - | Q1 | |
| [56] | STFA,SpSO | FS, HPT | DBN | - | 99.51 | - | Specificty=99.36, MCC=60.36 | Q1 | |
| [127] | TS | FS | EL | 13 | 99.5 | - | FAR=0.004 | Q1 | |
| N-BaIoT | [173] | BGWO | FS | EL (XGBoost) | - | 99.9941 | 99.9941 | - | Q2 |
| [103] | LOA-FOA | FS | RF | - | 99.86 | 99.86 | - | Q2 | |
| [132] | GWO-PSO | FS | RF | - | 99.86 | 99.86 | - | Q2 | |
| [17] | SSO-SA | FS | KNN | - | 98.7 | 99.8 | - | Q2 | |
| [18] | EPC | FS | KNN | - | 98.2 | 99.4 | - | Q1 | |
| [26] | APSO-WOA | HPT | CNN | N/A | 94.54 | - | JCC= 0.9 | Q1 |
Takeaway. According to the correlation analysis, the existing best IoT-IDSs have achieved 99.97%99.99% accuracy and 99.95%99.99% f1-score with NSL-KDD, UNSW-NB15, BoT-IoT, TON-IoT, and N-BaIoT datasets. Whereas, the performance slightly reduces while experimenting on the CICIDS-2017 and KDDCup-99 datasets (99.93%99.94% accuracy and f1-score). Overall, the metaheuristics and ML-integrated detection systems are effective in the IoT environment, irrespective of all widely used datasets. Another important observation is that existing intrusion detection datasets are extremely imbalanced since significant discrepancy is observed in data distribution between majority and minority classes (see Table VII). Consequently, it is not difficult to get a remarkable accuracy from any ML classifier [160], which can mislead researchers. Therefore, “accuracy” should not be considered a trustworthy performance metric in such scenarios. In contrast, the F1-score is widely accepted as a reliable metric since it reflects the harmonic balance between precision and recall [31]. For this reason, almost all top-notch IoT-IDSs measure F1-score along with accuracy and others.
@@
Further investigation in ensemble learning classifiers reveals the efficiency of using both traditional and deep machine learning algorithms. A significant portion of these classifiers are based on classic and deep ML, such as KNN+SVM+LSTM+MLP [67], RF+DT+MLP+KNN [145], and SVM+KNN+RF+LSTM [85]. In addition, a combination of classic ML classifiers is also observed, especially RFs [127], LR+RF+XGBoost [66], DT+AdaBoost+RF [45], and OC-SVM+IF+LOF [10]. Interestingly, Latif et al. [106] leveraged a CNN-based bootstrap ensemble classifier (Generic CNN+Xception+Inception+InseptionResntV2+EffcientNetV2L). Regarding optimization techniques, there is a consistent trend of employing nature-inspired metaheuristics, specifically FOA [152, 67], MFO [66], BGWO [173], CrSA [85], GWO [145], and GA [106]. In addition to these, search-based TS [127] along with hybrid metaheuristics BGSA-BGWO [45] and LS-PIO [10] are efficient in IoT-IDS. Moreover, almost all metaheuristics are used to select features; except in [152], where modified FOA tunes hyperparameters of the XGBoost classifier.
V-C2 Delve Into the Parameter and Hyperparameter Tuning Application